OpenStudy (anonymous):

find the lcd 1/7+1/4+1/3+1/6

OpenStudy (anonymous):

84 is the lcd

OpenStudy (anonymous):

1/7=12/84 1/4=21/84 1/3=28/84 1/6=14/84 12+21+28+14=65/84